OpenStudy (anonymous):

given v=5i-2j and w=-3i+j, find v * w

OpenStudy (goformit100):

use ther properties of complex number

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

\(v=5i-2j\) and \(w=-3i+j\) \[ v \times w=(5i-2j)\times(-3i+j)\]

OpenStudy (unklerhaukus):

wait, does * indicate the dot or cross product?

OpenStudy (anonymous):

HMMM

OpenStudy (chaise):

or does it indicate simple multiplication where you use FOIL?

OpenStudy (anonymous):

hey dork, is it dot or cross product?

OpenStudy (anonymous):

:)

OpenStudy (callisto):

@chaise Nope.. it looks like vectors

OpenStudy (anonymous):

yes its vectors

OpenStudy (callisto):

is it dot product or cross product (third time of this question..) ???

OpenStudy (anonymous):

dot?

OpenStudy (callisto):

v (dot) w =(5)(-3) + (-2)(1) = ... ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

im confused now xD

OpenStudy (callisto):

\(v = x_1i+y_1j\) and \(w= x_2i+y_2j\) v (dot) w = \(x_1x_2 + y_1y_2\)

OpenStudy (callisto):

given that they are 2 non-zero vectors

OpenStudy (unklerhaukus):

\[v=5i-2j=(-3,-2)\] \[w=-3i+j=(-3,1)\] \[v \cdot w=(5,-2)\cdot(-3,+1)\]\[\qquad=\left(5\times-3,\quad-2\times1\right)\]

OpenStudy (kropot72):

Convert the complex numbers to exponential form and use \[z _{1}z _{2}=r _{1}r _{2}e ^{i(\theta _{1}+\theta _{2})}\]

OpenStudy (unklerhaukus):

what complex numbers?

OpenStudy (unklerhaukus):

\[\{\hat i,\hat j\}\] are the unit vectors in the x, y dircetions

OpenStudy (kropot72):

Good question. j is used instead of i in electrical engineering.

OpenStudy (callisto):

Given two non-zero vectors \(v = x_1i+y_1j\) and \(w = x_2i+y_2j\) \[v \cdot w = (x_1i+y_1j) \cdot (x_2i+y_2j)\]\[v \cdot w = x_1i\cdot (x_2i+y_2j)+y_1j \cdot (x_2i+y_2j)\]\[v \cdot w = x_1x_2i\cdot i+x_1y_2i\cdot j+y_1x_2i\cdot j+y_1y_2j\cdot j\]\[v \cdot w = x_1x_2+0+0+y_1y_2\]\[v \cdot w = x_1x_2+y_1y_2\] Note that \(i \cdot i = 1\) and \(i\cdot j = 0\)